I have read and followed the above requirements fo Can someone help me with these NMR questions?
Answer #1

Double bond equivalent for

CaHbOcNdXe = a+1 - (b/2) - (d/2) + e/2

1.

Double bond equivalent for C5H9BrO2

C5H9BrO2 = 5 + 1 - (9/2) + (1/2) = 6 - 5 = 1

13C = 5 peaks

\Rightarrow 5 different types of carbon.

1H NMR

0.95 (t, 3H) - So, CH3 group and near a CH2 group present.

1.72 (m, 2H) - So, CH2 group and near more alkyl group present.

3.85 (s, 2H) - So, CH2 group with no adjacent carbon bearing hydrogen and Br group present (EWG)

4.15 (t, 2H) - So, CH2 group and near a CH2 group present. But it is at 4.15 ppm so, -COO group adjacent to it. The structure is shown in figure.

2.

C8H8O2 = 8 + 1 - (8/2) = 9 - 4 = 5

So, 5 double bond equivalent. 4 double bond and 1 ring

13C = 8 peaks

\Rightarrow 8 different types of carbon.

1H NMR

3.98 (s, 3H) - So, CH3 group no nearby no carbon bearing hydrogens.

7.60 (m, 4H) - So, 4 aromatic protrons.

10.00 (s, 1H) - So, aldehydic protron. The structure is shown in figure.

0 4.15 0.5 385 1.72 2. 760 7.60 าย 760 0